An air-plane has an effective wing surface area of 12.0 m² that is generating the lift force. In level flight the air speed over the top of the wings is 65.5 m/s, while the air speed beneath the wings is 39.0 m/s. What is the weight of the plane?(The density of air is 1.29 kg/m³)
Answer #1

The change in pressure is

P = \frac{1}{2}\rho \left ( v_2^2-v_1^2 \right )

\frac{F}{A} = \frac{1}{2}\rho \left ( v_2^2-v_1^2 \right )

F = \frac{A}{2}\rho \left ( v_2^2-v_1^1 \right )

F = \frac{12}{2}(1.29) \left ( (65.5)^2-(39)^2 \right )

F = 21434 N